Find two Laurent series in powers of z gor the function f defined by f(z)=1/z^2(1-z) and specify the regions in which the series converges to f(z).
1
Expert's answer
2020-05-26T19:51:40-0400

We remind the formula for geometric progression:

"\frac{1-z^{n+1}}{1-z}=1+z+z^2+\ldots +z^n,\quad z\neq1,\quad n\in{\mathbb{N}}." (1)

If "|z|<1" , we can take the limit "n\rightarrow\infty" and obtain

"\frac{1}{1-z}=1+z+z^2+z^3+\ldots =\sum\limits_{n=0}^{+\infty}z^n."

If we multiply (1) by "z^{-n}" , we get

"\frac{z^{-n}-z}{1-z}=z^{-n}+z^{1-n}+z^{2-n}+\ldots +1=1+z^{-1}+z^{-2}+\ldots+z^{-n},\quad n\in{\mathbb{N}}."

If "|z|>1" we take the limit "n\rightarrow\infty" and get

"\frac{z}{z-1}=1+z^{-1}+z^{-2}+z^{-3}+\ldots"

Thus, we obtain two Laurent series

"f(z)=\frac{1}{z^2(1-z)}=z^{-2}+z^{-1}+1+z+z^2+\ldots=\sum\limits_{n=-2}^{+\infty}z^n,\qquad 0<|z|<1,"

"f(z)=\frac{1}{z^2(1-z)}=-z^{-3}-z^{-4}-z^{-5}-\ldots=-\sum\limits_{n=-\infty}^{-3}z^n,\qquad |z|>1."